OpenStudy (anonymous):

30x-48y=33 20x+28y=37 solve by substitution method.

OpenStudy (anonymous):

pleasee help me.

OpenStudy (anonymous):

substitution for a system of equations says to get one variable in terms of the other and then substitute it into the second equation so that the only variable is the one you didn't solve for. so solve for x in the first equation, you get \[x = (33+48y)/30\] now plug into the second and you will get \[20(33+48y)/33 + 28y = 37 \] and doing some algebra gets y = 1/4 now go back and plug into the first equation and you get \[30x - 48(1/4) = 33\] and you get x = 3/2